A 12 Kg box is put on the surface of an inclined plane at 38 ° with the horizontal. The coefficient of kinetic friction is 0.35. Determine the magnitude of the force exerted by the inclined plane on the box.


1
Expert's answer
2021-10-24T18:24:38-0400

Given:

"m=12\:\rm kg"

"\theta=38^{\circ}"

"\mu=0.35"


The friction force

"F=\mu N=\mu mg\cos\theta"

"F=0.35*12*9.8*\cos38^{\circ}=32.4\:\rm N"